OpenStudy (anonymous):

write the equation in standard form of a parabola with a vertex of (-3,6) that goes through (1,-2).

OpenStudy (anonymous):

will somebody help me?

OpenStudy (anonymous):

Use the vertex form first; we know h=-3, k=6, x=1, y=-2; so we have: -2=a(1-(-3))²+6; solve it to get a=-1/2 so in vertex form it is: y = (-1/2)(x+3)²+6; expand and collect like terms to get the standard form y=(-1/2)x²-3x+3/2